# MultiFormer: A Multi-Person Pose Estimation System Based on CSI and Attention Mechanism

### v1

@@ II-B Feature Extraction of CSI
Recently, Transformer has become an important framework in the fields of natural language processing (NLP) and computer vision (CV). Researchers have turned to Transformer-based architecture due to its long-range dependency capture capability. For instance, Zhou et al. [22] introduced the CSI-Former model, which combines attention mechanisms with traditional CNNs to process CSI. This model allows the network to focus on those CSI subcarriers that are more important for pose estimation, thereby improving the efficiency and accuracy of feature extraction. Zhou et al. [20] proposed the Metafi++, which employs a CNN module to extract features from each pair of antennas, followed by a channel-wise self-attention mechanism to model interdependencies across different feature channels. Some studies inspired by Vision Transformers (ViT) have adopted the patch embedding idea [32][33] to analyze CSI spectra for human activity recognition.
@@ III-C1 Pose Probability Estimation
[…]
@@ IV-A Experimental Data Collection and Evaluation Metrics
Encoder 8 layers, Input sequence size 64$\times$1296 8 layers, Input sequence size 64$\times$576 6 layers, Input sequence size 64$\times$324
@@
IV-B Controlled Experiments
To demonstrate the advantages of the Time-Frequency Dual Token based Multi-Head Attention Network over single domain token (temporal or frequency) architectures in feature extraction, we compared the pose estimation performance of the MultiFormer under different token configurations. To ensure fair comparison, the single-token models share the same structure with the dual-token models by processing identical tokens (temporal or frequency) in both branches of the feature extractor.
TABLE III: Comparison of the PCK and Parameter size for Single and Dual Token MultiFormer Models on Collected Dataset
Feature Extractor PCK@5 PCK@10 PCK@20 Params (M)
Temporal Token 0.4697 0.7152 0.8813 11.93
Frequency Token 0.4702 0.7113 0.8502 11.93
Dual Token 0.5209 0.7501 0.8885 11.93
TABLE IV: PCK@5 Comparison of 18 Body Keypoints from Stage 1 to Stage 3 on Collected Dataset of MultiFormer
@@ IV-B Controlled Experiments
The performance of MultiFormer with different types of token is compared in Table III. As shown in Table III, the MultiFormer with the dual-token transformer architecture outperforms the models using either temporal or frequency tokens alone, despite maintaining identical parameter size (11.93M). This improvement shows the dual-token model’s capability to incorporate information from both temporal and frequency domains.
@@
We also conduct the experiment to show the effect of multi-stage pose estimation. In this test, Time-Frequency Dual Token-based Multi-Head Attention Network is used for feature extraction, and the estimation accuracy of different stages of MultiFormer is compared in Table IV. It can be seen from this example, as the number of stages increases, the PCK@5 of all keypoints improves. This enhancement is particularly notable for the keypoints with larger degrees of freedom that are challenging for the one-stage estimation to estimate. Specifically, compared to the one-stage estimation, the three-stage estimation improves the PCK@5 estimation accuracy by 12%–64% for all keypoints, with 47% for the right wrist and 64% for the left wrist. This shows that MSFN has successfully learned the interrelations and constraints among human pose keypoints.
IV-C Performance Comparison
To show the effectiveness of MultiFormer, we conduct a comprehensive comparison with several well-established models including WiSPPN[23], WPFormer[20], WiMose[25], CSI-Former[22], and CSI2Pose[27] on our self-collected dataset.
@@ IV-C Performance Comparison
TABLE V: PCK@5 Comparison on Collected Dataset for 18 Body Keypoints
Body Part MultiFormer CSI2Pose WiMose WPFormer CSIformer WISPPN
Nose 0.4634 0.2912 0.2900 0.2534 0.1869 0.1099
Neck 0.6069 0.4256 0.4237 0.3206 0.2848 0.1771
R-Shoulder 0.5845 0.4023 0.3754 0.3285 0.2599 0.1709
R-Elbow 0.4842 0.2910 0.2637 0.2118 0.1468 0.1083
R-Wrist 0.4268 0.2347 0.1787 0.1354 0.1080 0.0810
L-Shoulder 0.5797 0.3788 0.3686 0.2624 0.2526 0.1444
L-Elbow 0.4967 0.2939 0.2726 0.1656 0.1516 0.0958
L-Wrist 0.3853 0.1984 0.1830 0.1038 0.0823 0.0478
R-Hip 0.6151 0.4261 0.4042 0.2954 0.2300 0.1120
R-Knee 0.5533 0.3747 0.3382 0.1693 0.1669 0.1155
R-Ankle 0.5280 0.3324 0.2546 0.0519 0.1386 0.1146
L-Hip 0.6157 0.4301 0.3899 0.2621 0.2608 0.1238
L-Knee 0.5664 0.3813 0.3392 0.1327 0.1770 0.1381
L-Ankle 0.5369 0.3357 0.2538 0.1038 0.0998 0.1265
R-Eye 0.4680 0.2965 0.2929 0.2486 0.1994 0.1219
L-Eye 0.4712 0.3043 0.2971 0.2616 0.2015 0.1176
R-Ear 0.5001 0.3046 0.2882 0.2644 0.1986 0.1033
L-Ear 0.4945 0.3014 0.3024 0.2735 0.2167 0.0951
@@
TABLE VI: Comparison of the PCK@5, PCK@10, PCK@20 and Parameter size for Different Models on Collected Dataset
Method PCK@5 PCK@10 PCK@20 Params (M)
MultiFormer 0.5209 0.7501 0.8885 11.93
MultiFormer-24 0.3688 0.6765 0.8738 4.05
MultiFormer-18 0.2508 0.5561 0.8176 2.80
CSi2Pose 0.3335 0.6054 0.8175 17.90
WiMose 0.3064 0.6084 0.8418 76.65
WPFormer 0.2136 0.5328 0.8395 30.81
CSIFormer 0.1868 0.4755 0.7734 5.34
WISPPN 0.1169 0.3244 0.6390 20.19
@@
To give a detailed comparison between MultiFormer and other models, we show the PCK@5 comparison of the 18 skeletal keypoints in Table V for different pose estimation models on our self-collected dataset. In this test, MultiFormer achieves the highest estimation accuracies for all 18 keypoints, and the improvement of estimation accuracies for keypoints with high mobility is significant. For instance, MultiFormer improves the estimation accuracy for the right elbow from 0.2910 to 0.4842, which is more than 66.39% improvement, and greater improvement can be found for the estimation accuracies of left elbows and both wrists.
@@
Table VI shows the comparison of the average PCK@5, PCK@10, PCK@20 and parameter size between MultiFormer and its lightweight variants to other models on the self-collected dataset.
@@
Experiment results in Table VI demonstrate that MultiFormer achieves the highest estimation accuracy from PCK@5 to PCK@20. The estimation accuracy of the lightweight variant of MultiFormer, i.e., MultiFormer-24, achieves PCK@5 of 0.3688, which outperforms the tested model with similar parameter size. As we furthur decrease the parameter size of MultiFormer, the MultiFormer with a parameter size of 2.8M, i.e.,MultiFormer-18, still remains competitive performance.
@@
TABLE VII: Comparison of the PCK@20, PCK@30, PCK@40 and Parameter size for Different Models on MM-Fi Dataset
Method PCK@20 PCK@30 PCK@40 Params (M)
MultiFormer 0.7225 0.8065 0.8556 11.93
MultiFormer-24 0.6938 0.7925 0.8483 4.05
MultiFormer-18 0.6459 0.7643 0.8304 2.80
CSI2Pose 0.6841 0.7304 0.8139 17.90
HPE-Li 0.5207 0.6822 0.7818 1.66
WiMose 0.4867 0.6658 0.7734 36.20
WPFormer 0.4546 0.6444 0.7513 26.42
WISPPN 0.4541 0.6321 0.7408 26.78
@@
We further evaluate MultiFormer on the MM-Fi, a public benchmark dataset[35]. We compare the performance of MultiFormer and its lightweight variants on the MM-Fi[35] with several SOTA benchmarks including HPE-Li[37], WPFormer[20], and other models, i.e., WiMose[25], CSI2Pose[27], WiSPPN[23] in Table VII.
@@
As shown in Table VII, MultiFormer and its lightweight variants achieve better estimation accuracy on MM-Fi from PCK@20 to PCK@40 than other tested models. The lightweight variants of MultiFormer, i.e., MultiFormer-24(4.05M) and MultiFormer-18(2.80M) achieve average 0.6938 and 0.6459 for the PCK@20, respectively, which outperform other tested lightweight model.
@@
To compare the estimation accuracy for all 17 keypoints on MM-Fi between MultiFormer and other models with similar parameter sizes, we first compare the PCK@20 of MultiFormer(11.93M) with CSI2Pose(17.9M), which currently achieves the highest estimation accuracy among all tested models. The PCK@20 comparison of the 17 skeletal keypoints between MultiFormer and CSI2Pose is shown in Fig. 13. In this test, MultiFormer achieves higher estimation accuracy for all 17 keypoints.
@@
We also compare the PCK@20 of 17 keypoints on MM-Fi between lightweight models, i.e., MultiFormer-18(2.80M) and HPE-li(1.6M). The PCK@20 comparison of the 17 skeletal keypoints between MultiFormer-18 and HPE-Li is shown in Fig. 14. MultiFormer-18 outperforms HPE-li on the MM-Fi for 17 keypoints, particularly for high-mobility keypoints such as wrists and elbows. Specifically, the PCK@20 for the left and right hands are 0.0503 and 0.0508 for the HPE-Li [38], whereas our model achieves 0.3414 and 0.3334 with the improvement over 500%.
The tests shown in Fig. 13 and Fig. 14 together demonstrate that our proposed MultiFormer effectively captures the complex relationships and dependencies among human pose keypoints, thus, achieves better estimation accuracy.

### v2

@@ II-A CSI Preprocessing
Existing methods directly convert CSI into image-like tensors, which neglect the CSI’ temporal and frequency domains characteristics. Our proposed Time-Frequency Dual-Dimensional Tokenization (TFDDT) addresses this by creating time and frequency token streams, which enable more effective feature extraction.
@@ II-B Feature Extraction of CSI
Recently, Transformer has become an important framework in the fields of natural language processing (NLP) and computer vision (CV). Researchers have turned to Transformer-based architecture due to its long-range dependency capture capability. For instance, Zhou et al. [22] introduced the CSI-Former model, which combines attention mechanisms with traditional CNNs to process CSI. This model allows the network to focus on those CSI subcarriers that are more important for pose estimation, thereby improving the efficiency and accuracy of feature extraction. Zhou et al.
[20] proposed the Metafi++, which employs a CNN module to extract features from each pair of antennas, followed by a channel-wise self-attention mechanism to model interdependencies across different feature channels. Some studies inspired by Vision Transformers (ViT) have adopted the patch embedding idea [32][33] to analyze CSI spectra for human activity recognition.
Existing CNN-based feature extraction methods often struggle with capturing long-range dependencies and require significant computational resources. While Transformer-based models can capture long-range dependencies, they tend to neglect temporal dynamics. Our method’s dual-path attention captures both spectral and temporal features with the same or fewer parameter count and computational cost, giving better efficiency and performance.
@@ II-C Human Pose Estimation Techniques
Conventional single-stage pose estimation methods, i.e., keypoints regression, estimate the pose directly but fail to model the inter-keypoints connections properly. While the PAM-based models is able to capture the pair-wise links between keypoints, the full-body coherence is still a challeng- ing task. We address this task by cascading three heatmap- based feature fusion stages by Pose-Attentive modules that re- weight features spatially and channel-wise to give structurally consistent results.
@@ III-C1 Pose Probability Estimation
$W^{C}_{i}$ $\in\mathbb{R}^{1\times 256}=\text{PAPM}^{C}(PPH_{i-1})\quad\text{(Channel Attention)}$ (8)
$W^{S}_{i}$ $\in\mathbb{R}^{36\times 36}=\text{PAPM}^{S}(PPH_{i-1})\quad\text{(Spatial Attention)}$
@@ IV-A Experimental Data Collection and Evaluation Metrics
[…]
@@
We also use parameter count (model size) and FLOPs (Floating Point Operations Per Inference) to measure the computational efficiency of different models.
IV-B Ablation Study
IV-B1 Time–Frequency Dual-Tokenization
To evaluate the effect of the token design, we retrain the model by feeding only temporal tokens or only frequency tokens while keeping all other settings fixed. To ensure fair comparison, the single-token models share the same structure with the dual-token models by processing identical tokens (temporal or frequency) in both branches of the feature extractor.
TABLE III: Comparison of PCK, Parameter Count, and FLOPs for Single and Dual Token MultiFormer Models on Collected Dataset
Feature Extractor PCK@5 PCK@10 PCK@20 Params&FLOPs
Temporal Token 0.4697 0.7152 0.8813 11.93M&15.12G
Frequency Token 0.4702 0.7113 0.8502 11.93M&15.12G
Dual Token 0.5209 0.7501 0.8885 11.93M&15.12G
TABLE IV: PCK@5 Comparison of 18 Body Keypoints, Parameter Count, and FLOPs from Stage 1 to Stage 3 on Collected Dataset of MultiFormer
@@ IV-B1 Time–Frequency Dual-Tokenization
Params &FLOPs 1.95M &2.53G 6.85M &8.82G 11.93M &15.12G - - - -
The performance of MultiFormer with different token types is compared in Table III. As shown, the dual-token Transformer architecture improves the PCK@5 from 0.4697 (temporal-only) and 0.4702 (frequency-only) to 0.5209, i.e. an average improvement of roughly 11% , with the same parameter count of 11.93M. All three configurations also share the same computational cost of 15.12G FLOPs, confirming that the accuracy gain stems from dual-token modelling rather than extra computation. This result demonstrates the dual-token model’s ability to fuse information from both temporal and frequency domains more effectively than single-token counterparts.
@@
IV-B2 Three-Stage Feature-Fusion Decoder
We investigate the effectiveness of Multi-stage MSFN on the estiamtion accuracy by reducing the number of decoding stages. In this test, Time-Frequency Dual Token-based Multi-Head Attention Network is used for feature extraction, and the estimation accuracy of different stages of MultiFormer is compared in Table IV. It can be seen from this example, as the number of stages increases, the PCK@5 of all keypoints improves. This enhancement is particularly notable for the keypoints with larger degrees of freedom that are challenging for the one-stage estimation to estimate. Specifically, compared to the one-stage estimation, the three-stage estimation improves the PCK@5 estimation accuracy by 12%–64% for all keypoints, with 47% for the right wrist and 64% for the left wrist. This shows that MSFN has successfully learned the interrelations and constraints among human pose keypoints.
IV-B3 Channel-Spatial Attention in PAPM
This experiment measures how channel-level and spatial-level attention, individually and jointly, affect the pose-estimation quality.
TABLE V: Comparison of PCK, Parameter Count, and FLOPs for MultiFormer Models with different PAPM on Collected Dataset
Feature Extractor PCK@5 PCK@10 PCK@20 Params&FLOPs
Only Channel Attention 0.3520 0.6269 0.8292 11.93M&15.11G
Only Spatial Attention 0.3988 0.6569 0.8416 11.83M&15.12G
Full PAPM 0.5209 0.7501 0.8885 11.93M&15.12G
Table V shows that combining both channel-level and spatial-level attention lifts PCK@5 from 0.3520 (channel-only) and 0.3988 (spatial-only) to 0.5209, corresponding to relative improvements of 48.0% and 30.6%, respectively. Comparable gains appear for PCK@10 (up by 19.7% / 14.2%) and PCK@20 (up by 7.2% / 5.6%). All three models share an almost identical computational load (11.93 M parameters, 15.1G FLOPs), confirming that the accuracy boost comes from joint channel–spatial modelling rather than extra computation.
IV-C Performance Comparison With Existing Methods
To show the effectiveness of MultiFormer, we conduct a comprehensive comparison with several well-established models, including HPE-Li (2025)[37], CKDformer (2024)[38], WPFormer (2023)[20], CSI-Former (2022)[22], CSI2Pose (2021)[27], and WiSPPN (2019)[23] on our self-collected dataset.
@@ IV-C Performance Comparison With Existing Methods
TABLE VI: PCK@5 Comparison on Collected Dataset for 18 Body Keypoints
Body Part MultiFormer CSI2Pose HPE-Li CDKformer WPFormer
Nose 0.4634 0.2912 0.2760 0.2536 0.2534
Neck 0.6069 0.4256 0.3638 0.3838 0.3206
R-Shoulder 0.5845 0.4023 0.3419 0.3491 0.3285
R-Elbow 0.4842 0.2910 0.2604 0.2374 0.2118
R-Wrist 0.4268 0.2347 0.2016 0.1622 0.1354
L-Shoulder 0.5797 0.3788 0.3264 0.3306 0.2624
L-Elbow 0.4967 0.2939 0.2570 0.2385 0.1656
L-Wrist 0.3853 0.1984 0.1816 0.1631 0.1038
R-Hip 0.6151 0.4261 0.3772 0.3832 0.2954
R-Knee 0.5533 0.3747 0.3192 0.2960 0.1693
R-Ankle 0.5280 0.3324 0.2672 0.2259 0.0519
L-Hip 0.6157 0.4301 0.3766 0.3896 0.2621
L-Knee 0.5664 0.3813 0.3128 0.3095 0.1327
L-Ankle 0.5369 0.3357 0.2591 0.2365 0.1038
R-Eye 0.4680 0.2965 0.2801 0.2544 0.2486
L-Eye 0.4712 0.3043 02809 02603 0.2616
R-Ear 0.5001 0.3046 0.2793 02550 0.2644
L-Ear 0.4945 0.3014 0.2839 0.2805 0.2735
@@
TABLE VII: Comparison of the PCK@5, PCK@10, PCK@20, Parameter Count, and FLOPs for Different Models on Collected Dataset
Method PCK@5 PCK@10 PCK@20 Params&FLOPs
MultiFormer 0.5209 0.7501 0.8885 11.93M&15.12G
MultiFormer-24 0.3688 0.6765 0.8738 4.05M&2.24G
MultiFormer-18 0.2508 0.5561 0.8176 2.80M&1.10G
CSi2Pose[27] 0.3335 0.6054 0.8175 17.90M&32.41G
HPE-Li[37] 0.2914 0.5455 0.7833 1.66M&2.42G
CKDformer[38] 0.2783 0.5492 0.7993 33.52M&2.48G
WPFormer[20] 0.2136 0.5328 0.8395 30.81M&8.26G
CSIFormer[22] 0.1868 0.4755 0.7734 5.34M&6.24G
WISPPN[23] 0.1169 0.3244 0.6390 20.19M&7.70G
@@
To give a detailed comparison between MultiFormer and other models, we show the PCK@5 comparison of the 18 skeletal keypoints in Table VI for different pose estimation models on our self-collected dataset. In this test, MultiFormer achieves the highest estimation accuracies for all 18 keypoints, and the improvement of estimation accuracies for keypoints with high mobility is significant. For instance, MultiFormer improves the estimation accuracy for the right elbow from 0.2910 to 0.4842, which is more than 66.39% improvement, and greater improvement can be found for the estimation accuracies of left elbows and both wrists.
@@
Table VII shows the comparison of the average PCK@5, PCK@10, PCK@20 and parameter Count between MultiFormer and its lightweight variants to other models on the self-collected dataset.
@@
Experiment results in Table VII demonstrate that MultiFormer achieves the highest estimation accuracy from PCK@5 to PCK@20 while requiring only 15.12G FLOPs. The estimation accuracy of the lightweight variant of MultiFormer, i.e., MultiFormer-24, achieves PCK@5 of 0.3688 with just 2.24G FLOPs, which outperforms other tested models of comparable parameter size. Further reducing the model to 2.8 M parameters (MultiFormer-18) lowers the computational cost to 1.10G FLOPs yet still delivers competitive accuracy.
@@
TABLE VIII: Comparison of the PCK@20, PCK@30, PCK@40, Parameter Count, and FLOPs for Different Models on MM-Fi Dataset
Method PCK@20 PCK@30 PCK@40 Params&FLOPs
MultiFormer 0.7225 0.8065 0.8556 11.93M&15.12G
MultiFormer-24 0.6938 0.7925 0.8483 4.05M&2.24G
MultiFormer-18 0.6459 0.7643 0.8304 2.80M&1.10G
CSI2Pose[27] 0.6841 0.7304 0.8139 17.90M&32.41G
DT-Pose[39] 0.6580 0.7790 0.8510 36.51M&2.70G
CKDformer[38] 0.5219 0.6843 0.7813 33.52M&2.48G
HPE-Li[37] 0.5207 0.6822 0.7818 1.66M&2.42G
WPFormer[20] 0.4546 0.6444 0.7513 26.42M&507.89G
WISPPN[23] 0.4541 0.6321 0.7408 26.78M&159.81G
@@
We further evaluate MultiFormer on the MM-Fi, a public benchmark dataset[35]. We compare the performance of MultiFormer and its lightweight variants on the MM-Fi[35] with several SOTA benchmarks including HPE-Li (2025)[37], DT-Pose (2025)[39], CKDformer (2024)[38], WPFormer (2023)[20], CSI2Pose (2021)[27], and WiSPPN (2019)[23] in Table VIII.
@@
As shown in Table VIII, MultiFormer and its lightweight variants achieve better estimation accuracy on MM-Fi from PCK@20 to PCK@40 than other tested models. The lightweight variants of MultiFormer, i.e., MultiFormer-24(4.05M) and MultiFormer-18(2.80M) achieve average 0.6938 and 0.6459 for the PCK@20, respectively, which outperform other tested lightweight model.
@@
To compare the estimation accuracy for all 17 keypoints on MM-Fi between MultiFormer and other models with similar parameter sizes, we first compare the PCK@20 of MultiFormer (11.93 M / 15.12G FLOPs) with CSI2Pose (17.9 M / 32.41G FLOPs), which currently achieves the highest estimation accuracy among all tested models. The PCK@20 comparison of the 17 skeletal keypoints between MultiFormer and CSI2Pose is shown in Fig. 13. In this test, MultiFormer achieves higher estimation accuracy for all 17 keypoints while using less than half the computation.
@@
We also compare the PCK@20 of 17 keypoints on MM-Fi between lightweight models, i.e., MultiFormer-18(2.80M) and HPE-li(1.6M).
The PCK@20 comparison of the 17 skeletal keypoints between MultiFormer-18 (2.80 M / 1.10G FLOPs) and HPE-Li (1.6 M / 2.42G FLOPs) is shown in Fig. 14. MultiFormer-18 outperforms HPE-li on the MM-Fi for 17 keypoints with roughly 55 % less computation, particularly for high-mobility keypoints such as wrists and elbows. Specifically, the PCK@20 for the left and right hands are 0.0503 and 0.0508 for the HPE-Li [37], whereas our model achieves 0.3414 and 0.3334 with the improvement over 500%.